# Bayesian RG Flow in Neural Network Field Theories

### v1

@@ 1 Introduction
A quintessential method for creating flows in the space of SFTs is via Renormalization Group (RG) schemes [^3] Here, we are broadly defining RG schemes to encompass both perturbative and non-perturbative approaches. While the latter is more general, we note that plainly saying “RG” colloquially refers to the former. Additional terms (i.e. “Wilsonian”, “Exact”, “functional”, etc.) are often prepended to refer to the latter. To avoid confusion, in this work, we will exclusively be considering a very general non-perturbative Exact RG (ERG) scheme – Bayesian RG (BRG), which holds Wilsonian and Polchinski’s ERG as special cases.. In the picture proposed by Wilson [35, 36], and refined by many others [28], SFTs can be connected via ERG equations which describe how the SFT changes with scale. Within physics contexts, many traditional SFTs have the property of spatial locality, thus establishing a natural scale with which to view these flows: energy (or equivalently, distance). A traditional SFT viewed at low-energies (large distances) will generally be insensitive to high-energy (small distance) effects. [^4] More precisely, the high-energy effects can be captured by modifying the couplings between fields at low energies. An ERG flow connects a traditional SFT at high-energies (UV theory) to its corresponding low-energy counterpart (IR theory) by systematically aggregating small-distance effects, i.e. coarse-graining. However, the SFT duals to NN architectures are most often spatially non-local; therefore, the ERG-like coarse graining of these NN models, and, by extension, their dual SFTs, must be organized according to a non-spatial scale. To this end, the authors in Refs. [30, 31] have introduced a more general scale which lends itself to spatially non-local settings: information-theoretic similarity. In particular, they showed that ERG-like coarse-graining of NN parameters with respect to this scale amounts to inverting the process of Bayesian Inference; thus this procedure has been dubbed Bayesian RG (BRG). On the other hand, recall that training a BNN is a step-by-step application of Bayesian Inference. Utilizing the NNFT correspondence, we can interpret training as inducing an inverse BRG flow in SFT space from the information-theoretic ‘IR’ to a ‘UV’ fixed-point determined by the training data and objective. In certain circumstances, the BRG method becomes equivalent to traditional ERG schemes where the scale is based on spatial locality, e.g. Polchinski’s ERG [37].
@@
Fig. 1: Here, $\pi(\theta)$ is the parameter distribution of a NN $\phi_{\theta}$ which has a dual SFT with action, $S[\phi]$. The BRG scheme coarse grains the original NN parameter distribution to $\pi_{\Lambda}$, indexed by Fisher information scale, $\Lambda$. Updated NN ensembles have a transformed dual action, $S_{\Lambda}[\phi]$, connected to original action through an information-theoretic BRG flow.
@@
This unified approach provides a pathway to present NN architectures in terms of dual SFTs, and the reverse. Establishing this connection could have many potentially interesting applications at the interface of physics and machine learning. For example, this correspondence potentially provides a method to construct representations of exotic SFTs so that they can be accurately and efficiently sampled. We note that this proposed application is highly related, although technically distinct, from existing literature [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49] leveraging methods such as diffusion models [50, 51] or normalizing flows [52] for the same tasks (see Section 5 for a more detailed discussion). As another example, this correspondence could be used to systematically explore the space of NNFTs and also potentially used as a new framework to study interesting phenomena in NN training behavior (e.g. criticality [20, 21, 22, 25], grokking [53], feature learning [54, 55, 56, 57, 58, 59, 60], scaling laws [61, 62, 63, 64]).
@@
The current work sets the stage for these and other applications. In Section 2, we begin by reviewing Bayesian Inference with the training of NN ensembles as a special case. Section 3 establishes the connection between the NNFT and BRG frameworks. In Section 3.1, we first review the NNFT correspondence as introduced in Refs. [32, 33, 34, 12] and use this to connect the pair $(\phi_{\theta},\pi)$, representing a NN architecture and its parameter distribution, to its dual SFT with action, $S(\phi)$. Then, in Section 3.2, we review the BRG scheme [30, 31] for arbitrary inference models and discuss the similarities and differences between BRG and conventional Wilsonian ERG schemes. In Section 3.3, we then apply the BRG technique directly to NN ensembles. For a fixed architecture, BRG defines a family of parameter distributions indexed by the distinguishability scale, $\Lambda$. We then theoretically show that these parameter distributions can be passed through the NNFT correspondence to produce a family of SFTs with actions, $\{S_{\Lambda}(\phi)\}$, connected through an information-theoretic BRG flow. In Section 4, we analytically and numerically demonstrate the BRG-NNFT correspondence on analytically tractable cases. In Section 4.1, we explicitly compute the BRG flow for trained NNs with infinite width and arbitrary depth and activation functions. We find formulae for the renormalized mean ($1$-pt function) and covariance ($2$-pt function) which parameterize the NN as a renormalized Gaussian Process in this limit. In Section 4.1.1, we specialize these results to the case of a single layer NN with the generalized cos-net activation considered previously in Refs. [32, 33]. This allows us to take advantage of the fact that, with a well chosen initialized $\pi(\theta)$, this architecture corresponds to a free scalar SFT having spectrum identical to conventional free scalar SFT. In this case, we show that BRG explicitly renormalizes the mass of the SFT, and can be interpreted in terms of a more familiar contraction of the momentum space of the SFT. Finally, in Section 4.2, we implement this joint scheme in a series of numerical experiments which corroborate our analytic results for asymptotically-wide NNs with ReLU [65] activations. We conclude with a discussion of our results in Section 5 where we also provide recommendations for future directions.
@@ 1.1 Related works
In this work, we will be using the NNFT correspondence described in Refs. [32, 33, 34, 12], which views a NN output function as a parameterization of a field, $\phi$, sampled from an SFT with action, $S[\phi]$. The NN itself is a mapping from $d$-dimensional Euclidean space to an output: $x\rightarrow\phi_{\theta}(x)$ with parameters $\theta$ and inputs $x\in\mathbb{R}^{d}$. However, there are many other works [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] which use alternate formulations to connect NNs with SFTs. For example, Refs. [9, 20] define a series of fields at each pre-activation, allowing them to make statements about the inter-layer interactions of these fields. There is also a rich literature, using methods from statistical physics to theoretically study many aspects of NN behavior [16, 66, 17]. There is also a closely related body of literature studying the theoretical behavior of Bayesian NNs (BNNs) [67, 68, 59, 69, 56, 70, 71, 72, 73]. As we will describe in subsequent sections, the present work connects directly to the training of BNNs. In particular, BRG-NNFT inherits the same computational difficulties as BNNs in estimating posterior distributions [74]. In this work, we consider a sample-estimate of posterior distributions by considering an ensemble of NNs. However, it would be interesting to explore BRG-NNFT in the context of other posterior estimation schemes [74, 75, 76, 1, 77, 78] in future work.
@@
Many previous works [79, 80, 81, 82, 83, 84, 26, 27, 85, 86, 9, 20, 14, 39, 30, 31, 87] have also connected aspects of NNs to various RG schemes. In particular, various RG schemes have been used to study criticality and inter-layer information propagation [86, 9, 20], to understand NN generalization [14], and also to establish a direct relationship between Polchinski’s ERG and the training of diffusion models [39, 30, 31, 87]. Furthermore, machine learning methods utilizing an information-theoretic RG scheme based on Real Space Mutual Information (RSMI) have also been used as a way to construct physically relevant operators from data [80, 81, 82, 83]. While these works also utilize a related information-theoretic quantity (Mutual Information), the relevance of operators are determined based on their long-distance behavior i.e. the cutoff scale is still based on a notion of spatial locality. Another series of works [26, 27] claims that the training of an ensemble of NNs can be interpreted as inducing an RG flow in the corresponding SFTs. To demonstrate this, they incrementally changed the weight distribution of their NN ensemble as a proxy for training, and showed that this induces an RG flow in the corresponding SFTs. In this work, we go a step further to actually train the NN ensemble explicitly, and then subsequently renormalize with BRG. Additionally, while both Refs. [26, 27] and our present analysis focus on the infinite-width GP limit of networks, we stress that the BRG-NNFT correspondence holds for any architecture. Our consideration of the infinite-width GP limit is precisely to demonstrate the BRG-NNFT correspondence in an analytically tractable case that makes contact with existing results.
@@ 2 Preliminaries
In this section we provide an introduction to Bayesian inference for the purpose of performing and understanding machine learning. This section does not intend to provide a technical review of strategies for implementing Bayes [^6] In exposition we will use the abbreviation ‘Bayes’ to refer to Bayesian inference. in the context of machine learning [^7] For those interested in such a survey, see e.g. [3]., but rather a dictionary translating the mathematical tools and ideas from Bayes into those that are more familiar to readers of the machine learning literature. We begin by introducing Bayesian inference in the most general context of a parameterized probability model, and conclude by addressing how machine learning for NNs can be treated as a special case of such an approach. The familiar reader can safely skip this section and move to Section 3. For the reader who is in a rush, a summary of this section can be found in Table 1.
@@ 2.1 Introduction to Bayesian Inference
Bayesian inference principally consists of two stages, model selection and learning. The model selection stage begins by understanding the structure of the system one is interested in learning about. In particular, we let $(S,d^{d}y)$ be a measure space in which a random variable $Y$ takes its value. The stochastic properties of the variable $Y$ are encoded by specifying a probability distribution over $S$. A probability distribution is a measurable, normalized function $p\in L^{1}(S,d^{d}y)$ e.g.
@@
Returning to the problem of statistical inference, to implement Bayes’ law we should regard $(D,\Theta)$ as defining a joint random variable on the measure space $S^{\times|D|}\times M$ [^9] The measure on this space is a tensor product of $|D|$ copies of the measure on $S$ with the measure on $M$ e.g. $d^{d}y_{1}\;\cdots d^{d}y_{|D|}\;d^{n}\theta\;$ Here we have regarded the sample $D=\{y_{1},...,y_{|D|}\}$ as a collection of independent and identically distributed random variables taking values in $S$. The distribution for $D$ conditional on $\theta$ is given by
@@
As the notation suggests, the posterior predictive can be interpreted as the “expected value” over the model class (6) with respect to the weighting defined by the posterior. In summary, Bayesian inference can be regarded as the arrow
@@ 3 NNFT + BRG
Fig. 2: A commutative diagram which unifies NNFT and BRG. The composition of these maps defines a new proposal for constructing information geometric flows through the space of field theories. We interpret these BRG flows as a suitable replacement for standard field theoretic ERG flows in general contexts where spatially local coarse graining is not sufficient for implementing a meaningful flow.
Each of the three subsections of this section coincide with one leg of the diagram (Figure 2). In Section 3.1 we introduce the Neural Network Field Theory (NNFT) correspondence, which is a tool for translating network architectures with stochastic parameters into explicit statistical field theories (SFTs). In Section 3.2 we establish the set up for Bayesian Renormalization, which is an information geometry inspired approach to coarse graining over the space of models in a statistical inference experiment. Finally, in Section 3.3 we combine NNFT and BRG to obtain an information theoretic renormalization group for exploring the space of SFTs.
@@ 3.1 Neural Network-Field Theory Correspondence
Fig. 3: NNFT maps a fixed NN architecture, $\phi_{\theta}$, with parameter distribution, $\pi(\theta)$, into a Euclidean field theory with action, $S[\phi]$.
@@
$\mathbb{E}_{p}(F)=\int D\phi\int_{P}d^{n}\theta\;\pi(\theta)\;\delta(\phi-\phi_{\theta})F(\phi)=\int_{P}d^{n}\theta\;\pi(\theta)\;F(\phi_{\theta})=E_{\pi}(F(\phi_{\theta})).$ (28)
@@
In [32] it has been shown that $p(\phi)\propto e^{-S(\phi)}$ has the form of a generically interacting Euclidean field theory with action $S(\phi)=S_{\text{free}}(\phi)+S_{\text{int}}(\phi)$ given by
@@
In (34), $g^{(n)}(x_{1},...,x_{n})$ are space dependent couplings obtained from the connected correlation functions (cumulants) via a set of ‘inverse Feynman rules’ [32]. The free action is well-defined for models with an invertible, i.e. has full rank, $2$-pt function, with $G^{(2)}_{c}(x,y)^{-1}$ regarded as the inverse of the two point function or the network kernel. Thus, there is a direct correspondence between a generic NN and an SFT with Euclidean action $S(\phi)$. This correspondence, introduced in Refs. [12, 34], is summarized in Table 2.

### v2

@@ Abstract
Copyright attribution to authors.
This work is a submission to SciPost Physics Core.
License information to appear upon publication.
Publication information to appear upon publication.
Received Date
Accepted Date
Published Date
@@ 1 Introduction
A quintessential method for creating flows in the space of SFTs is via renormalization group (RG) schemes. [^3] Here, we are broadly defining RG schemes to encompass both perturbative and non-perturbative approaches. While the latter is more general, we note that plainly saying “RG” colloquially refers to the former. Additional terms (i.e. “Wilsonian”, “Exact”, “functional”, etc.) are often prepended to refer to the latter. To avoid confusion, in this work, we will exclusively be considering a very general, non-perturbative exact RG (ERG) scheme – Bayesian RG (BRG), which holds Wilsonian and Polchinski’s ERG as special cases. In the picture proposed by Wilson [35, 36], and refined by many others [28], SFTs can be connected via ERG equations which describe how the SFT changes with scale. Within physics contexts, many traditional SFTs have the property of spatial locality, thus establishing a natural scale with which to view these flows: energy (or, equivalently, distance). A traditional SFT viewed at low-energies (large distances) will generally be insensitive to high-energy (small distance) effects. [^4] More precisely, the high-energy effects can be captured by modifying the couplings between fields at low energies. An ERG flow connects a traditional SFT at high-energies (UV theory) to its corresponding low-energy counterpart (IR theory) by systematically aggregating small-distance effects, i.e. coarse-graining. However, the SFT duals to NN architectures are most often spatially non-local; therefore, the ERG-like coarse graining of these NN models, and, by extension, their dual SFTs, must be organized according to a non-spatial scale. To this end, the authors in Refs. [30, 31] have introduced a more general scale which lends itself to spatially non-local settings: information-theoretic similarity. In particular, they showed that ERG-like coarse-graining of NN parameters with respect to this scale amounts to inverting the process of Bayesian inference; thus, this procedure has been dubbed Bayesian RG (BRG). On the other hand, recall that training a BNN is a step-by-step application of Bayesian inference. Utilizing the NNFT correspondence, we can interpret training as inducing an inverse BRG flow in SFT space from the information-theoretic ‘IR’ to a ‘UV’ fixed-point determined by the training data and objective. In certain circumstances, the BRG method becomes equivalent to traditional ERG schemes where the scale is based on spatial locality, e.g. Polchinski’s ERG [37].
@@
Figure 1: Here, $\pi(\theta)$ is the parameter distribution of a NN, $\phi_{\theta}$, which has a dual SFT with action, $S[\phi]$. The BRG scheme coarse grains the original NN parameter distribution to $\pi_{\Lambda}$, indexed by the Fisher information scale, $\Lambda$. Updated NN ensembles have a transformed dual action, $S_{\Lambda}[\phi]$, connected to original action through an information-theoretic BRG flow.
@@
This unified approach provides a pathway to present NN architectures in terms of dual SFTs, and the reverse. Establishing this connection could have many potentially interesting applications at the interface of physics and machine learning. For example, this correspondence potentially provides a method to construct representations of exotic SFTs so that they can be accurately and efficiently sampled. We note that this proposed application is highly related, although technically distinct, from existing literature [39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49] leveraging methods such as diffusion models [50, 51] or normalizing flows [52] for the same tasks (see Section 5 for a more detailed discussion). As another example, this correspondence could be used to systematically explore the space of NNFTs and also potentially used as a new framework to study interesting phenomena in NN training behavior (e.g. criticality [20, 21, 22, 25], grokking [53], feature learning [54, 55, 56, 57, 58, 59, 60, 61, 62], scaling laws [63, 64, 65, 66, 61, 67]).
@@
The current work sets the stage for these and other applications. In Section 2, we begin by reviewing Bayesian inference with the training of NN ensembles as a special case. Section 3 establishes the connection between the NNFT and BRG frameworks. In Section 3.1, we first review the NNFT correspondence as introduced in Refs. [32, 33, 34, 12] and use this to connect the pair $(\phi_{\theta},\pi)$, representing a NN architecture and its parameter distribution, to its dual SFT with action, $S[\phi]$. Then, in Section 3.2, we review the BRG scheme [30, 31] for arbitrary inference models and discuss the similarities and differences between BRG and conventional Wilsonian ERG schemes. In Section 3.3, we then apply the BRG technique directly to NN ensembles. For a fixed architecture, BRG defines a family of parameter distributions indexed by the distinguishability scale, $\Lambda$. We then theoretically show that these parameter distributions can be passed through the NNFT correspondence to produce a family of SFTs with actions, $\{S_{\Lambda}[\phi]\}$, connected through an information-theoretic BRG flow. In Section 4, we analytically and numerically demonstrate the BRG-NNFT correspondence on analytically tractable cases. In Section 4.1, we explicitly compute the BRG flow for trained NNs with infinite width and arbitrary depth and activation functions. We find formulae for the renormalized mean ($1$-pt function) and covariance ($2$-pt function) which parameterize the NN as a renormalized Gaussian Process in this limit. In Section 4.1.1, we specialize these results to the case of a single layer NN with the generalized cos-net activation considered previously in Refs. [32, 33]. This allows us to take advantage of the fact that, with a well-chosen initialized $\pi(\theta)$, this architecture corresponds to a free scalar SFT having spectrum identical to conventional free scalar SFT. In this case, we show that BRG explicitly renormalizes the mass of the SFT, and can be interpreted in terms of a more familiar contraction of the momentum space of the SFT. Finally, in Section 4.2, we implement this joint scheme in a series of numerical experiments which corroborate our analytic results for asymptotically-wide NNs with ReLU [68] activations. We conclude with a discussion of our results in Section 5 where we also provide recommendations for future directions.
@@ 1.1 Related works
In this work, we will be using the NNFT correspondence described in Refs. [32, 33, 34, 12], which views a NN output function as a parameterization of a field, $\phi$, sampled from an SFT with action, $S[\phi]$. The NN itself is a mapping from $d$-dimensional Euclidean space to an output: $x\rightarrow\phi_{\theta}(x)$ with parameters $\theta$ and inputs $x\in\mathbb{R}^{d}$. However, there are many other works [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19] which use alternate formulations to connect NNs with SFTs. For example, Refs. [9, 20] define a series of fields at each pre-activation, allowing them to make statements about the inter-layer interactions of these fields. There is also a rich literature, using methods from statistical physics to theoretically study many aspects of NN behavior [16, 69, 17, 61, 62, 67], as well as a closely related body of literature studying the theoretical behavior of Bayesian NNs (BNNs) [70, 71, 59, 72, 56, 73, 74, 75, 76]. As we will describe in subsequent sections, the present work connects directly to the training of BNNs. In particular, BRG-NNFT inherits the same computational difficulties as BNNs in estimating posterior distributions [77]. In this work, we consider a sample-estimate of posterior distributions by considering an ensemble of NNs. However, it would be interesting to explore BRG-NNFT in the context of other posterior-estimation schemes [77, 78, 79, 1, 2, 80] in future work.
@@
Many previous works [81, 82, 83, 84, 85, 86, 26, 27, 87, 88, 9, 20, 14, 39, 30, 31, 89] have also connected aspects of NNs to various RG schemes. In particular, various RG schemes have been used to study criticality and inter-layer information propagation [88, 9, 20], to understand NN generalization [14], and also to establish a direct relationship between Polchinski’s ERG and the training of diffusion models [39, 30, 31, 89]. Furthermore, machine learning methods utilizing an information-theoretic RG scheme based on real space mutual information (RSMI) have also been used as a way to construct physically relevant operators from data [82, 83, 84, 85]. While these works also utilize a related information-theoretic quantity (mutual information), the relevance of operators are determined based on their long-distance behavior i.e. the cutoff scale is still based on a notion of spatial locality. Another series of works [26, 27] claims that the training of an ensemble of NNs can be interpreted as inducing an RG flow in the corresponding SFTs. To demonstrate this, they incrementally changed the weight distribution of their NN ensemble as a proxy for training, and showed that this induces an RG flow in the corresponding SFTs. In this work, we go a step further to actually train the NN ensemble explicitly, and then subsequently renormalize with BRG. Additionally, while both Refs. [26, 27] and our present analysis focus on the infinite-width GP limit of networks, we stress that the BRG-NNFT correspondence holds for any architecture. Our consideration of the infinite-width GP limit is precisely to demonstrate the BRG-NNFT correspondence in an analytically-tractable case that makes contact with existing results.
@@ 2 Preliminaries
In this section we provide an introduction to Bayesian inference for the purpose of performing and understanding machine learning. This section does not intend to provide a technical review of strategies for implementing Bayes [^6] In exposition, we will use the abbreviation ‘Bayes’ to refer to Bayesian inference. in the context of machine learning [^7] For those interested in such a survey, see e.g. [3]., but rather a dictionary translating the mathematical tools and ideas from Bayes into those that are more familiar to readers of the machine learning literature. We begin by introducing Bayesian inference in the most general context of a parameterized probability model, and conclude by addressing how machine learning for NNs can be treated as a special case of such an approach. The familiar reader can safely skip this section and move to Section 3. For the reader who is in a rush, a summary of this section can be found in Table 1.
@@ 2.1 Introduction to Bayesian Inference
Bayesian inference principally consists of two stages, model selection and learning. The model selection stage begins by understanding the structure of the system one is interested in learning about. In particular, we let $(S,d^{d}y)$ be a measure space in which a random variable, $Y$, takes its value. The stochastic properties of the variable $Y$ are encoded by specifying a probability distribution over $S$. A probability distribution is a measurable, normalized function $p\in L^{1}(S,d^{d}y)$ e.g.
@@
Returning to the problem of statistical inference, to implement Bayes’ law we should regard $(D,\Theta)$ as defining a joint random variable on the measure space $S^{\times|D|}\times M$ [^9] The measure on this space is a tensor product of $|D|$ copies of the measure on $S$ with the measure on $M$ e.g. $d^{d}y_{1}\;\cdots d^{d}y_{|D|}\;d^{n}\theta\;$. Here we have regarded the sample $D=\{y_{1},...,y_{|D|}\}$ as a collection of independent and identically distributed random variables taking values in $S$. The distribution for $D$ conditional on $\theta$ is given by
@@
As the notation suggests, the posterior predictive can be interpreted as the “expected value" over the model class (6) with respect to the weighting defined by the posterior. In summary, Bayesian inference can be regarded as the arrow
@@ 3 NNFT + BRG
Figure 2: A commutative diagram which unifies NNFT and BRG. The composition of these maps defines a new proposal for constructing information geometric flows through the space of field theories. We interpret these BRG flows as a suitable replacement for standard field theoretic ERG flows in general contexts where spatially local coarse graining is not sufficient for implementing a meaningful flow.
Each of the three subsections of this section coincide with one leg of the diagram (Figure 2). In Section 3.1, we introduce the Neural Network Field Theory (NNFT) correspondence, which is a tool for translating network architectures with stochastic parameters into explicit statistical field theories (SFTs). In Section 3.2, we establish the set up for Bayesian Renormalization, which is an information geometry inspired approach to coarse graining over the space of models in a statistical inference experiment. Finally, in Section 3.3, we combine NNFT and BRG to obtain an information theoretic renormalization group for exploring the space of SFTs.
@@ 3.1 Neural Network-Field Theory Correspondence
Figure 3: NNFT maps a fixed NN architecture, $\phi_{\theta}$, with parameter distribution, $\pi(\theta)$, into a Euclidean field theory with action, $S[\phi]$.
@@
$\mathbb{E}_{p}(F)=\int D\phi\int_{P}d^{n}\theta\;\pi(\theta)\;\delta(\phi-\phi_{\theta})F(\phi)=\int_{P}d^{n}\theta\;\pi(\theta)\;F(\phi_{\theta})=\mathbb{E}_{\pi}(F(\phi_{\theta})).$ (28)
@@
In [32] it has been shown that $p(\phi)\propto e^{-S[\phi]}$ has the form of a generically interacting Euclidean field theory with action $S[\phi]=S_{\text{free}}(\phi)+S_{\text{int}}(\phi)$ given by
@@
In (34), $g^{(n)}(x_{1},...,x_{n})$ are space dependent couplings obtained from the connected correlation functions (cumulants) via a set of ‘inverse Feynman rules’ [32]. The free action is well-defined for models with an invertible, i.e. has full rank, $2$-pt function, with $G^{(2)}_{c}(x,y)^{-1}$ regarded as the inverse of the two point function or the network kernel. Thus, there is a direct correspondence between a generic NN and an SFT with Euclidean action $S[\phi]$. This correspondence, introduced in Refs. [12, 34], is summarized in Table 2.